### v2

# LLM-as-a-Judge: Reassessing the Performance of LLMs in Extractive QA

Abstract
Extractive reading comprehension question answering (QA) datasets are typically evaluated using Exact Match (EM) and F1-score, but these metrics often fail to fully capture model performance. With the success of large language models (LLMs), they have been employed in various tasks, including serving as judges (LLM-as-a-judge). In this paper, we reassess the performance of QA models using LLM-as-a-judge across four reading comprehension QA datasets. We examine different families of LLMs and various answer types to evaluate the effectiveness of LLM-as-a-judge in these tasks. Our results show that LLM-as-a-judge is highly correlated with human judgments and can replace traditional EM/F1 metrics. By using LLM-as-a-judge, the correlation with human judgments improves significantly, from 0.22 (EM) and 0.40 (F1-score) to 0.85. These findings confirm that EM and F1 metrics underestimate the true performance of the QA models. While LLM-as-a-judge is not perfect for more difficult answer types (e.g., job), it still outperforms EM/F1, and we observe no bias issues, such as self-preference, when the same model is used for both the QA and judgment tasks. [^1] Our data and code are available at https://github.com/Alab-NII/llm-judge-extract-qa
1 Introduction
Machine reading comprehension (MRC) is a crucial task for evaluating natural language understanding, designed to test a model’s reading comprehension by requiring it to answer questions based on a given text (Hirschman et al., 1999). Many datasets have been proposed over the past several years, such as SQuAD (Rajpurkar et al., 2016; 2018) for simple MRC, QuAC (Choi et al., 2018) and CoQA (Reddy et al., 2019) for conversational MRC, and QAngaroo (Welbl et al., 2018), HotpotQA (Yang et al., 2018), and FanOutQA (Zhu et al., 2024) for multi-hop MRC.
Based on the answer format, Chen (2018) classify existing MRC tasks into four types: extraction, multiple-choice (MC), cloze-style, and free-form. Some recent datasets also introduce yes/no answers (Clark et al., 2019; Geva et al., 2021). Depending on the answer type, corresponding evaluation metrics are used. Since the yes/no, MC, and cloze-style (selecting an answer from a set of options) typically do not face issues with evaluation metrics–accuracy being the standard–other datasets may encounter difficulties. For instance, datasets with extractive answer types often rely on Exact Match (EM) and F1-score as evaluation metrics. However, we observe that these two metrics do not accurately reflect the models’ actual performance, often underestimating it, as shown in Figure 1. The core issue stems from the fact that while there are many valid ways to phrase a correct answer, the evaluation process usually relies on a single, fixed gold answer. This lack of flexibility in accepting alternate yet correct phrasings leads to an underestimation of the model’s true performance. Another issue is that generative AI models vary in style; some generate complete sentences containing the answer, while others may repeat part of the question before answering.
Figure 1: An example that shows EM and F1-score underestimate the performance of models, while LLM-as-a-judge provides a more robust perspective.
Recently, the success and capabilities of large language models (LLMs) have led to their application in various tasks, including their use as judges (LLM-as-a-judge) (Zheng et al., 2023; Chiang & Lee, 2023; Kocmi & Federmann, 2023; Li et al., 2024; Verga et al., 2024; Chen et al., 2024). LLM-as-a-judge is used for tasks involving text generation, such as machine translation (Kocmi & Federmann, 2023), text summarization (Liu et al., 2023; Skopek et al., 2023; Wu et al., 2023), story generation (Chiang & Lee, 2023), and reading comprehension QA (Zheng et al., 2023; Verga et al., 2024; Chen et al., 2024). Verga et al. (2024) is the most similar to ours. They experiment with different QA datasets to demonstrate that, rather than using a single large model as a judge, we can combine several smaller models to form juries. Our work complements this study by delving deeper into the reading comprehension QA task to reveal the true capabilities of current LLMs. Unlike their approach, we provide fine-grained analyses to examine the impact of LLM-as-a-judge on specific answer types. Moreover, we will release our dataset, which can be used to evaluate the use of LLM-as-a-judge in the extractive QA task, as they did not release theirs.
In this paper, we aim to reassess the performance of LLMs on the reading comprehension QA task by employing LLM-as-a-judge. To gain a more comprehensive understanding of the current capabilities of LLMs in reading comprehension QA tasks, we conduct experiments on four different datasets with multiple types of answers: Quoref (Dasigi et al., 2019), DROP (Dua et al., 2019), HotpotQA (Yang et al., 2018), and 2WikiMultiHopQA (Ho et al., 2020). Additionally, to gain deeper insights into the effectiveness of LLM-as-a-judge for evaluating reading comprehension QA tasks, we analyze the existing answer types to identify where LLM-as-a-judge is beneficial and where it falls short. Furthermore, we use different families of LLMs for both the QA and LLM-as-a-judge tasks, and we also conduct an analysis of self-preference bias (Liu et al., 2024; Panickssery et al., 2024) when the same model is used for both tasks.
Our results show that using LLM-as-a-judge is highly correlated with human judgments, improving from 0.22 (EM) and 0.40 (F1-score) to 0.85, highlighting its potential to replace EM and F1 metrics. This high correlation further confirms that EM and F1 metrics underestimate the true performance of the models. Our analysis demonstrates that LLMs as judges are particularly effective for answer types related to numbers and dates in extractive QA. Moreover, we observe no self-preference bias in the extractive QA task when the same model is used for both the QA and judgment tasks.
2 Related Work
Extractive QA.
An extractive QA task requires a model to extract a span of text from the provided context to answer a given question. Early datasets for this task include SQuAD (Rajpurkar et al., 2016; 2018), CNN/DailyMail (Hermann et al., 2015), and NewsQA (Trischler et al., 2017). Later datasets introduced additional challenges, such as multi-hop reasoning (Welbl et al., 2018; Yang et al., 2018; Trivedi et al., 2022), coreferential reasoning (Dasigi et al., 2019), and numerical reasoning (Dua et al., 2019). To the best of our knowledge, previous extractive QA datasets have predominantly relied on automatic evaluation metrics, such as EM, F1, or accuracy, in their default evaluation protocols.
Evaluation of Extractive QA.
In addition to evaluating a predicted answer for the main question task using EM and F1, some studies also consider using EM and F1 for evaluating explanation tasks (Yang et al., 2018; Ho et al., 2020), such as sentence-level supporting facts prediction in HotpotQA (Yang et al., 2018) or evidence prediction in 2WikiMultiHopQA (Ho et al., 2020) and $\mathrm{R^{4}C}$ (Inoue et al., 2020). Other studies evaluate sub-questions as explanations for the answers but often rely on EM, F1, accuracy, or a compositionality gap derived from accuracy for their assessments (Tang et al., 2021; Press et al., 2023; Wu et al., 2025).
LLM-as-a-Judge.
Thanks to their success and capabilities, LLMs have been applied to a variety of tasks, including serving as judges for tasks related to text generation, such as text summarization (Liu et al., 2023; Skopek et al., 2023; Wu et al., 2023) and QA (Zheng et al., 2023; Verga et al., 2024; Chen et al., 2024). However, several issues arise when using LLMs as judges, particularly concerning their accuracy and reliability in this role. One natural approach is to compare the scores given by LLM-as-a-judge with human scores to assess their agreement or correlation (Thakur et al., 2025). Previous work has highlighted several biases in LLM-as-a-judge, such as order bias (Wang et al., 2024) and egocentric bias (Koo et al., 2024). Another concern is that LLMs can be vulnerable to adversarial attacks, which could manipulate their scoring (Shi et al., 2024; Raina et al., 2024). For a more detailed discussion on the use of LLM-as-a-judge, we refer readers to the comprehensive survey paper by Gu et al. (2025).
In this paper, we focus on using LLM-as-a-judge for the extractive QA task, which is simpler than other text generation tasks. We employ various LLMs as judges in our experiments and compare their evaluations with human judgments to assess the correlation scores of LLM-as-a-judge in this context.
3 Experimental Settings
3.1 Datasets
We select four reading comprehension QA datasets for our experiments based on the following criteria: (1) they use EM and F1 as evaluation metrics, (2) they have not yet been fully solved, (3) they provide context for the given questions, and (4) they contain diverse types of answers. These four datasets are: Quoref (Dasigi et al., 2019), DROP (Dua et al., 2019), HotpotQA (Yang et al., 2018), and 2WikiMultiHopQA (2Wiki; Ho et al., 2020). To ensure consistency in our comparisons, we only use datasets that employ EM and F1 as evaluation metrics. To assess whether the current EM and F1 scores may underestimate the true performance of the models, we select datasets where these scores are not excessively high (e.g., below 95%). To simplify the judgement process, we choose datasets that provide the corresponding context, allowing us to refer to it when determining whether a predicted answer is acceptable. Regarding answer types, we select datasets from various tasks, such as numerical reasoning and multi-hop reasoning, to ensure a diverse range of answer types for our analysis.
Quoref.
Quoref is a reading comprehension dataset designed to assess the coreferential reasoning abilities of models. In the default dataset setup, a passage can contain multiple QA pairs. We treat each pair as an individual sample in our experiments. The questions are created through crowdsourcing, with a focus on the coreference resolution phenomenon. Most of the answer types are in string format, such as person names.
DROP.
DROP is a reading comprehension dataset designed to evaluate the numerical reasoning abilities of models. Similar to Quoref, a single paragraph can contain multiple QA pairs, and we treat each pair as an individual sample. It is worth noting that, in addition to extractive answer types, DROP also includes number and date answer types. These two types may not always be extractive (i.e., a span of text appearing directly in the context), such as when a question asks about the next day of a specific day. However, we include them in our analyses, as they do not pose challenges for generative models in predicting the correct answer.
HotpotQA.
HotpotQA is a multi-hop QA task that requires multiple reasoning steps to answer each question. The dataset comprises two main question types: bridge and comparison. Notably, the presence of comparison questions introduces yes/no answers into the dataset. HotpotQA is designed for two key tasks: answer prediction and sentence-level supporting facts prediction. In our work, we focus exclusively on the answer prediction task.
2Wiki.
Similar to HotpotQA, 2Wiki also features two main question types, bridge and comparison. In addition, it includes a separate task designed to evaluate the explanatory abilities of models. As with HotpotQA, our focus is solely on the answer prediction task.
Obtaining Answer Types.
From the development set of each dataset, we use heuristic rules to obtain the answer type. We define the rules for 8 answer types: string, place, name, job, date, number, year, and bool as follows:
  • String: questions that cannot be categorized into the following categories.
  • Place: questions starting with “where” or asking about the city, country, state, place, region, nationality, etc.
  • Name: questions starting with “who” and “whom” or asking about the name, nickname, role, player, actor, etc.
  • Job: questions asking about occupation, job, profession, or career.
  • Date: questions asking about which or what date.
  • Number: questions starting with “how many”, “how much”, or asking about percentage, population, etc.
  • Year: questions asking about which or what year.
  • Bool: yes-no questions.
Our goal is to carefully select a representative subset of approximately 1,000 samples from each dataset to ensure sufficient and meaningful analysis. It is important to note that our rules may not capture every question from the full development set. To ensure diversity in answer types for our experiment, we include all samples from answer types with fewer than 100 instances. For answer types exceeding this threshold, we randomly sample in proportion to their representation in the entire dataset. Since the boolean answer type (yes/no) does not pose significant challenges when evaluated using EM and F1 metrics, we exclude samples with boolean answers from our experiments. Table 1 presents the number of samples for each answer type, along with the total number of samples for each dataset used in our experiments.
Dataset String Place Name Job Date Number Year Total
Quoref 55 74 878 23 - 21 - 1,051
DROP 124 24 107 5 18 740 - 1,018
HotpotQA 146 257 460 92 78 77 14 1,124
2Wiki 296 383 292 - 29 - - 1,000
Total 621 738 1,737 120 125 838 14 4,193
Table 1: The number of samples per answer type and the total number of samples for the four datasets. A dash (‘-’) indicates that the corresponding answer type is not present in the dataset. In the case of Quoref, we observed that the date and year answer types were frequently mispredicted by rules due to the specific nature of the related questions. Therefore, we chose to exclude these samples from our experiments.
3.2 Models
QA Task.
We use four different model families for the QA task: Mistral v0.1 (7B and 8x7B) (Jiang et al., 2023; 2024), Qwen 2 (7B and 72B) (Yang et al., 2024), Gemma 2 (9B and 27B) (Team, 2024), and Llama 3.1 (8B and 70B) (Grattafiori et al., 2024).
LLM-as-a-Judge.
We use three model families as LLM-as-a-judge systems: Mistral-Instruct-7B-v0.3 (Jiang et al., 2023), Llama 3.3 70B (Grattafiori et al., 2024) and Qwen 2.5 72B (Yang et al., 2025). It is noted that these model families are similar to those used in our QA task, but we employ an enhanced version. To examine the self-preference bias (Liu et al., 2024; Panickssery et al., 2024), we also run Qwen 2 (7B and 72B) and Llama 3.1 (8B and 70B) as judges in our analyses. Notably, all the models used in our experiments are instruction-tuned.
3.3 Promptings
Since the development of in-context learning (Brown et al., 2020), more advanced prompting techniques have been introduced, such as chain-of-thought (CoT) prompting (Wei et al., 2022) and zero-shot CoT prompting (Kojima et al., 2022).
QA Task.
Since most current LLMs are familiar with the QA task and the datasets we use often involve lengthy context passages, coupled with our focus on evaluation, we chose to adopt zero-shot CoT prompting in our experiments. Our prompt is presented in Appendix A.1.
LLM-as-a-Judge.
To evaluate the predicted answers using LLMs, we use a few-shot prompting approach. In this approach, we provide a few demonstration examples to guide the model in labeling the answers. Due to length constraints and for the sake of simplicity, these demonstrations exclude the context. During evaluation, we present the model with a question, a gold answer, a predicted answer, and the corresponding context. The model is then instructed to assign one of two labels to the predicted answer: Correct and Incorrect.
  • CORRECT: The predicted answer matches the gold answer or is a valid alternative (e.g., different but correct ways of writing a name).
  • INCORRECT: The predicted answer is wrong or does not align with the gold answer.
We use a 6-shot approach in our judgment prompt. We randomly select five samples from each dataset (the subset not used in our evaluation) and run a simple QA model to generate predicted answers. Subsequently, we manually choose four shots from these samples and reuse two exemplars from the paper by Verga et al. (2024). The prompt used in our study is presented in Appendix A.1.
4 Results and Analyses
4.1 Reliability of LLM-as-a-Judge Scores
To assess the reliability of LLM-as-a-judge scores, we first collect human judgments on the predicted answers. We then calculate the correlation between these human judgments and the LLM-as-a-judge scores.
Human Judgements.
We sampled 200 instances (50 per dataset), maintaining the original distribution of answer types. Each sample included all 8 predicted answers. Annotators (two authors of this paper) were presented with the question, gold answer, predicted answers, and context, and asked to label each predicted answer as either ‘Correct or ‘Incorrect’. Since the task is simple, we decided to assign only one annotator to each sample. However, annotators may discuss ambiguous cases. We excluded cases where the gold answer was incorrect, leaving 161 valid samples, each with 8 predicted answers, resulting in 1,288 predicted answers. These were used to calculate the correlation between human judgment and LLM performance as a judge.
Correlation to Human Judgements.
We calculate the Pearson correlation (Pearson, 1895) between human judgments and the LLM-as-a-judge evaluations from three models: Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B. The correlation scores are presented in Table 2. As shown in the table, Qwen 2.5 exhibits the highest correlation with human judgments, followed by Llama 3.3, while Mistral v0.3 has the lowest correlation scores. According to standard explanation, a correlation score above 0.80 is generally considered strong. We focus on using Qwen 2.5 72B as the judge for our subsequent analyses.
Additionally, we include the correlation scores between human judgments and the EM/F1 scores. For the F1-score, we use a threshold of 0.5 to classify values between 0 and 1 for correlation calculation. As presented in the table, these two metrics, EM/F1 scores, have correlation scores smaller than those of all LLM-as-a-judge models. These scores indicate that we can use LLM-as-a-judge to evaluate the extractive QA task.
QA Model Mistral 7B Llama 3.3 70B Qwen 2.5 72B EM F1
Mistral 7B v0.1 0.627 0.627 0.851 0.157 0.311
Mixtral 8x7B 0.592 0.549 0.723 0.126 0.244
Qwen 2 7B 0.598 0.781 0.863 0.234 0.434
Qwen 2 72B 0.653 0.680 0.793 0.157 0.274
Gemma 2 9B 0.715 0.833 0.848 0.288 0.572
Gemma 2 27B 0.711 0.850 0.908 0.247 0.487
Llama 3.1 8B 0.700 0.876 0.945 0.331 0.628
Llama 3.1 70B 0.626 0.803 0.848 NaN 0.281
Average 0.653 0.750 0.847 0.220 0.404
Table 2: Pearson correlation coefficients between human judgments and the LLM-as-a-judge evaluations from three models—Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B—across eight different QA models are presented. Additionally, we include the correlation scores between human judgments and the EM/F1 scores. NaN values arise when none of the predicted answers by Llama 3.1 70B exactly match the gold answers, resulting in a constant list of 0s, for which correlation is undefined. This often occurs when gold answers are plain numbers (e.g., 93.5), but predictions include extra context like percentages (e.g., 93.5%) or time spans (e.g., 38 years).
4.2 Comparing EM/F1 Scores and LLM-as-a-Judge
We first present the performance gap between EM/F1 scores and LLM-as-a-judge for false samples – those with a false EM score. Next, we compare EM/F1 scores and LLM-as-a-judge across all samples. Ideally, LLM-as-a-judge should be used only when the EM score is false.
Using LLM-as-a-Judge for False Samples.
Table 3 presents the EM and F1 scores, along with the LLM-as-a-judge scores (from three models: Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B) for the four datasets: Quoref, DROP, HotpotQA, and 2Wiki. As shown in the Table, the LLM-as-a-judge scores from the three models are higher than both the EM and F1 scores. The largest gap is 81.9 between EM and Llama 3.3 70B as a judge on HotpotQA, evaluating Mixtral 8x7B’s answers. The smallest gap is 15.9 between EM and Qwen 2.5 72B as a judge on 2Wiki, evaluating Gemma 2 9B’s answers. However, this comparison may be biased since LLM-as-a-judge is only used for false samples, while true EM scores are directly assigned to it. To offer a broader perspective, we present experimental results for all samples in the following section.
Model Quoref DROP
EM F1 Mistral Llama Qwen EM F1 Mistral Llama Qwen
Mistral 7B v0.1 1.2 14.9 65.6 80.5 57.9 1.6 11.3 59.8 49.4 40.9
Mixtral 8x7B 13.9 31.1 85.6 86.5 81.2 0.1 8.8 69.2 72.4 60.6
Qwen 2 7B 33.7 46.6 76.0 64.2 61.1 18.1 27.1 65.2 54.4 48.5
Qwen 2 72B 45.5 62.3 91.8 88.8 87.1 13.9 25.6 82.1 84.6 78.3
Gemma 2 9B 53.2 68.8 86.0 82.3 78.1 40.4 49.2 77.6 74.8 70.0
Gemma 2 27B 58.6 72.7 88.4 82.6 80.5 45.6 56.0 80.4 80.1 75.3
Llama 3.1 8B 46.3 61.9 83.0 73.0 70.7 34.0 45.1 68.5 62.8 58.8
Llama 3.1 70B 61.0 76.9 93.8 90.6 89.7 34.2 50.8 87.3 87.4 83.3
HotpotQA 2WikiMultihopQA
EM F1 Mistral Llama Qwen EM F1 Mistral Llama Qwen
Mistral 7B v0.1 14.8 32.2 85.9 80.9 75.5 10.1 26.5 65.7 56.7 46.7
Mixtral 8x7B 7.1 26.1 85.9 89.0 84.6 6.9 25.3 73.1 73.7 66.4
Qwen 2 7B 41.2 56.9 90.8 86.3 82.4 36.4 47.9 77.3 65.9 60.4
Qwen 2 72B 38.4 55.0 94.6 94.4 91.5 33.1 48.0 84.7 85.3 78.5
Gemma 2 9B 56.4 72.9 94.0 91.9 88.7 52.7 61.3 77.2 77.9 68.6
Gemma 2 27B 56.6 74.3 94.7 93.1 89.9 53.1 63.8 78.9 79.5 71.8
Llama 3.1 8B 51.8 68.2 91.5 89.1 85.9 45.7 55.1 80.8 71.2 63.1
Llama 3.1 70B 54.0 71.2 95.8 94.9 92.8 59.3 68.3 85.6 87.2 81.9
Table 3: Automatic evaluation scores (EM and F1) and LLM-as-a-judge scores (highlighted in green and blue) from three models – Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B – for the four datasets.
Using LLM-as-a-Judge for All Samples.
Table 10 (in Appendix B.1) presents LLM-as-a-judge scores for Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B across four datasets, evaluated in two cases: false EM samples only and all samples. As shown in the table, the gaps between these evaluations are minimal, indicating comparable performance.
Summary.
Table 3 highlights the performance gap between EM/F1 scores and LLM-as-a-judge scores. Our correlation analysis between human judgments and Qwen-as-a-judge confirms that EM and F1 scores underestimate model performance – not because LLM-as-a-judge incorrectly labels predictions as correct, but because these metrics fail to capture true performance. Notably, using LLM-as-a-judge for false EM samples only or all samples yields similar results. In practice, we apply LLM-as-a-judge to false EM samples only.
4.3 Analyses
Type #Samples Correlation
name 464 0.862
number 336 0.899
place 240 0.771
string 160 0.862
job 72 0.352
date 16 1.000
Figure 2: Correlation scores of Qwen-as-a-judge with human judgment for each answer type.
Figure 3: Percentage of each answer type identified as correct by Qwen-as-a-judge.
Effects of LLM-as-a-Judge on Different Answer Types.
We use 1,288 predicted answers with human judgment in Section 4.1 and LLM-as-a-judge scores from Qwen on these samples, and calculate the correlation between the two for each answer type. The scores are presented in Figure 3. As expected, the answer types “date” and “number” show the highest correlation score, while the scores for other types are relatively similar. We observe that for the answer type “job”, the correlation score is quite low, primarily due to the ambiguity of multiple jobs in the gold answer, as the predicted answer can contain more or fewer jobs, making the judgment more difficult. Table 4.3 presents several examples highlighting these differences. We also provide additional examples for other answer types.
Example Question Gold Answer Predicted Answer Type Labels
1 What profession did Willi Forst and Elmer Clifton share? actor Actor, screenwriter, film director, film producer and singer Job Human: False Model: True
2 What is Michael Nakasone job at the college prep school in Honolulu? teaches and conducts Michael Nakasone is a band director at Punahou School Job Human: False Model: True
3 What was the profession of the one who wrote a song on a 2005 album he collaborated on with Marc Predka? professional wrestler, actor, and hip hop musician hip hop artist Job Human: False Model: True
4 What date did the take over of Enniscorthy end? 30 April The takeover of Enniscorthy … exact date is not provided in the context. However, since the takeover started on Thursday, 27 April, and lasted until Sunday, we can infer that the takeover ended on 30 April 1916. Date Human: True Model: True
5 Why did the director of film The Obsessed Of Catule die? stroke Anselmo Duarte died due to complications from a stroke String Human: True Model: True
6 How many field goals between 25 and 40 yards were made? 2 Two field goals between 25 and 40 yards were made Number Human: True Model: True
7 Who was involved in observing the natural progression of untreated syphilis in rural African-American men in Alabama? John Charles Cutler The U.S. Public Health Service Name Name: False Model: False
8 Where was the place of burial of Amun-Her-Khepeshef’s mother? QV66 The provided text does not mention where Amun-Her-Khepeshef’s mother was buried. Place Name: False Model: False
Table 4: Examples of differences and similarities between labels provided by human judgment and LLM-as-a-judge (with labels from Qwen used in this case).
In Figure 3, we show the percentage of each answer type that LLM-as-a-judge identified as correct when the EM score is false. As shown in the figure, the judge is able to correctly predict more than 50% of each answer type, except for “date”. The highest scores are achieved for the “job” and “place” answer types. However, when considering the correlation score with human judgment for the “job” answer type, we observe that LLM-as-a-judge is less strict than humans when judging the correctness of the predicted answer regarding jobs.
Self-Preference.
We consider four QA models (Llama 3.1 8B, Llama 3.1 70B, Qwen 2 7B, and Qwen 2 72B) and seven LLM-as-a-judge models (Llama 3.1 8B, Llama 3.1 70B, Llama 3.3 70B, Qwen 2 7B, Qwen 2 72B, Qwen 2.5 72B, and Mistral 7B). Suppose we have seven judgment models, denoted as $m_{1}$ to $m_{7}$ and the QA model is $m_{1}$. We define self-preference bias as the scenario in which all remaining models unanimously label the predicted answer as incorrect (threshold = 100%), while only $m_{1}$ judges its own answer as correct. We also calculate this score for different thresholds, such as 83%, where one of the six remaining models can judge the answer as correct. To measure this bias, we calculate the percentage of such cases relative to the total number of cases where the EM score is false. Table 6 presents an example of self-preference bias where the QA model is the same as Judge 1.
Table 6 presents the percentage of self-preference bias scores for the four models across three different thresholds. These scores indicate a relatively small self-preference bias for Llama 3.1 8B when it serves as both a QA model and a judge model. However, for the other models, the percentages are much smaller. This can be explained by the fact that, in the extractive QA task, where the gold answer is clearly provided, it is more difficult for self-bias to occur compared to other tasks.
Judge 1 CORRECT
Judge 2 INCORRECT
Judge 3 INCORRECT
Judge 4 INCORRECT
Judge 5 INCORRECT
Judge 6 INCORRECT
Judge 7 INCORRECT
Table 5: Example of self-preference bias where the QA model is the same as Judge 1.
QA Model Threshold = 100% 83% 67%
Llama 3.1 8B 5.77 12.04 14.77
Llama 3.1 70B 0.26 0.77 1.62
Qwen 2 7B 0.63 2.06 4.48
Qwen 2 72B 0.14 0.34 0.85
Table 6: Percentage of self-preference bias scores for the four models across three different thresholds.
We present examples of self-preference bias for different QA models in Table 4.3. We observe that in these cases, the predicted answer is obviously wrong, and all remaining models label it as incorrect. However, only the judge model, which is the same as the QA model, is predicted as correct.
Example Question Gold Answer Predicted Answer QA Model Labels
1 What is Abigail’s nickname? Abby Abigail Llama 3.1 8B judge 1: CORRECT judge 2: INCORRECT judge 3: INCORRECT judge 4: INCORRECT judge 5: INCORRECT judge 6: INCORRECT judge 7: INCORRECT
2 What was the name of the person that Fauré taught? Ravel Ralph Vaughan Williams Llama 3.1 8B judge 1: CORRECT judge 2: INCORRECT judge 3: INCORRECT judge 4: INCORRECT judge 5: INCORRECT judge 6: INCORRECT judge 7: INCORRECT
3 Where did Mei Shaowu’s father die? Peking The context does not provide information about Mei Shaowu’s father’s death location Llama 3.1 70B judge 1: INCORRECT judge 2: CORRECT judge 3: INCORRECT judge 4: INCORRECT judge 5: INCORRECT judge 6: INCORRECT judge 7: INCORRECT
4 What is the last name of the person who is friends with Cap’n Billau? O’Conner Billau Qwen 2 7B judge 1: INCORRECT judge 2: INCORRECT judge 3: INCORRECT judge 4: CORRECT judge 5: INCORRECT judge 6: INCORRECT judge 7: INCORRECT
5 How many yards longer was the longest field goal compared to the shortest? 31 18 yards Qwen 2 72B judge 1: INCORRECT judge 2: INCORRECT judge 3: INCORRECT judge 4: INCORRECT judge 5: CORRECT judge 6: INCORRECT judge 7: INCORRECT
Table 7: Examples of self-preference bias for different QA models.
5 Conclusion
In this paper, we investigate the use of LLM-as-a-judge to reassess the performance of QA models. With generative AI models, predicted answers for extractive QA tasks can take various forms that traditional metrics like EM and F1-score fail to capture. We find that LLMs as judges provide a more comprehensive evaluation, revealing that EM and F1-score underestimate the true capabilities of the models. Our results show that LLM-as-a-judge is highly correlated with human judgments and can effectively replace traditional EM/F1 metrics. Although LLM-as-a-judge is not perfect for more difficult answer types (e.g., job), it still outperforms EM/F1, and we observe no bias issues, such as self-preference, when the same model is used for both the QA and judgment tasks.
Appendix A Experimental Settings
A.1 Promptings
Table 8 presents the QA task prompt that we used in our experiment.
System Prompt: You are an expert in question answering systems.
User Prompt: Answer the following question based on the provided context.
### Instructions:
* Task: Identify the correct answer from the provided context.
* Approach:
- Break down the problem into smaller parts, if necessary.
- Carefully reason through your answer step-by-step.
- Ensure that your answer is directly supported by the context.
### Response Format:
Please format your answer within brackets as follows:
  <ans> Your Answer </ans>
* Question: {question}
* Context: {context}
Table 8: The QA task prompt.
Table 9 presents the LLM-as-a-judge task prompt used in our experiments.
System Prompt: You are an expert in question answering systems.
User Prompt: Your job is to evaluate a predicted answer by comparing it against the gold answer and the given question.
You may refer to the provided context if needed.
## Grading Criteria:
* CORRECT: The predicted answer matches the gold answer or is a valid alternative (e.g., different but correct ways of writing a name).
* INCORRECT: The predicted answer is wrong or does not align with the gold answer.
* In some ambiguous cases, where it is unclear whether the predicted answer is correct or not, please refer to the provided context and use it as the final source for making your judgment.
## Response Format:
Please format your answer within brackets as follows: <ans> Your Answer </ans> ## Here are some examples:
### Example 1:
* Question: What nationality is the performer of song Daddy, Come Home?
* Gold Answer: United States
* Predicted Answer: American
* Label: <ans> CORRECT </ans>
### Example 2:
* Question: Who is Bohemond Iv Of Antioch’s paternal grandmother?
* Gold Answer: Constance of Antioch
* Predicted Answer: princess Constance of Antioch
* Label: <ans> CORRECT </ans>
### Example 3:
* Question: Rejuvelac is kind of grain water invented and promoted by a ’holistic health’ practitioner born in which year?
* Gold Answer: 1909
* Predicted Answer: Rejuvelac is a kind of grain water invented and promoted by Ann Wigmore, who was born in 1909.
* Label: <ans> CORRECT </ans>
### Example 4:
* Question: What is the birthday of the actress who was the Duchess in ’The Revengers Tragedy’?
* Gold Answer: 23 November 1946
* Predicted Answer: Diana Quick, who played the Duchess in ’The Revengers Tragedy’, was born on 23rd September 1934.
* Label: <ans> INCORRECT </ans>
### Example 5:
* Question: Who beacme a star as a comic book character created by Gerry Conway and Bob Oksner?
* Gold Answer: Megalyn Echikunwoke
* Predicted Answer: Vixen
* Label: <ans> INCORRECT </ans>
### Example 6:
* Question: How many years after the peace treaty was annulled did the Sauk finally grant the Fox sanctuary?
* Gold Answer: 3
* Predicted Answer: 11 years
* Label: <ans> INCORRECT </ans>
## Here is your task:
* Question: {question}
* Gold Answer: {gold_ans}
* Predicted Answer: {pred_ans}
* Context: {context}
Table 9: The LLM-as-a-judge task prompt.
Appendix B Results and Analyses
B.1 Comparing EM/F1 Scores and LLM-as-a-Judge
Table 10 presents LLM-as-a-judge scores for Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B across four datasets. Columns labeled Mistral, Llama, or Qwen show scores for false EM samples only, while Mistral-A, Llama-A, or Qwen-A show scores for all samples.
Model Quoref DROP
Mistral Mistral-A Llama Llama-A Qwen Qwen-A Mistral Mistral-A Llama Llama-A Qwen Qwen-A
Mistral 7B v0.1 65.6 65.7 80.5 80.6 57.9 57.6 59.8 60.2 49.4 49.8 40.9 41.0
Mixtral 8x7B 85.6 86.0 86.5 87.0 81.2 82.3 69.2 72.8 72.4 75.9 60.6 62.7
Qwen 2 7B 76.0 75.9 64.2 64.2 61.1 60.8 65.2 65.2 54.4 54.4 48.5 48.0
Qwen 2 72B 91.8 91.7 88.8 88.8 87.1 87.1 82.1 82.1 84.6 84.6 78.3 78.1
Gemma 2 9B 86.0 85.6 82.3 82.3 78.1 77.8 77.6 77.6 74.8 74.8 70.0 70.1
Gemma 2 27B 88.4 88.0 82.6 82.6 80.5 80.4 80.4 80.4 80.1 80.1 75.3 75.6
Llama 3.1 8B 83.0 82.7 73.0 72.9 70.7 70.6 68.5 68.5 62.8 62.8 58.8 58.9
Llama 3.1 70B 93.8 93.6 90.6 90.6 89.7 89.3 87.3 87.3 87.4 87.4 83.3 83.2
HotpotQA 2WikiMultihopQA
Mistral Mistral-A Llama Llama-A Qwen Qwen-A Mistral Mistral-A Llama Llama-A Qwen Qwen-A
Mistral 7B v0.1 85.9 85.9 80.9 80.8 75.5 75.1 65.7 65.6 56.7 56.8 46.7 46.5
Mixtral 8x7B 85.9 88.5 89.0 91.5 84.6 87.0 73.1 73.6 73.7 74.3 66.4 66.5
Qwen 2 7B 90.8 90.8 86.3 86.2 82.4 82.4 77.3 77.0 65.9 65.8 60.4 60.8
Qwen 2 72B 94.6 94.4 94.4 94.3 91.5 92.0 84.7 84.3 85.3 85.2 78.5 78.7
Gemma 2 9B 94.0 93.8 91.9 91.8 88.7 88.3 77.2 76.6 77.9 77.8 68.6 69.0
Gemma 2 27B 94.7 94.6 93.1 93.0 89.9 89.4 78.9 78.6 79.5 79.5 71.8 72.0
Llama 3.1 8B 91.5 91.4 89.1 88.7 85.9 85.8 80.8 80.5 71.2 71.2 63.1 63.2
Llama 3.1 70B 95.8 95.6 94.9 94.8 92.8 92.4 85.6 85.1 87.2 87.2 81.9 81.9
Table 10: LLM-as-a-judge scores for Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B across four datasets. Columns labeled Mistral, Llama, or Qwen indicate scores on false EM samples only, while Mistral-A, Llama-A, or Qwen-A indicate scores on all samples.